1. A ray of light from the sun strikes the windshield of a car at an angle of incidence of 36.0° and refracts at an angle of 24.0°. If the initial light had a wavelength of 630 nm, determine the wavelength of the light after it has refracted.  
1
Expert's answer
2020-12-10T11:06:46-0500
"n=\frac{\sin{36}}{\sin{24}}\\n\lambda'=\lambda\\
\\\frac{\sin{36}}{\sin{24}}\lambda'=630\\\lambda'=436\ nm"